### v1

# A structure theorem for rooted connectivity in bidirected graphs

Abstract.
Recently, bidirected graphs have received increasing attention from the graph theory community with both structural and algorithmic results. Bidirected graphs are a generalization of directed graphs, consisting of an undirected graph together with a map assigning each endpoint of every edge either sign ++ or -. The connectivity properties of bidirected graphs are more complex than those of directed graphs and not yet well understood.
In this paper, we show a structure theorem about rooted connectivity in bidirected graphs in terms of directed graphs. As applications, we prove Lovász’ flame theorem, Pym’s theorem and a strong variant of Menger’s theorem for a class of bidirected graphs and provide counterexamples in the general case.
Key words and phrases:
Menger’s Theorem, flames, Pym’s theorem, bidirected graph, connectivity
2020 Mathematics Subject Classification:
05C40 (Primary) 05C38, 05C20 (Secondary)
T.A. is supported by the National Science Foundation Award Number DMS-2303251 and the Alexander von Humboldt Foundation. A.J. was funded by the Deutsche Forschungsgemeinschaft (DFG) - Grant No. 513023562. F.R. is supported by a doctoral scholarship of the Studienstiftung des deutschen Volkes. Q.T. is supported by China Scholarship Council (No. 202206770031) and Sino-German (CSC-DAAD) Postdoc Scholarship Program, 2022 (57607866).
1. Introduction
A bidirected graph is a graph $B$ endowed with a sign function $\sigma$ that assigns a sign $\{+,-\}$ to each end of every edge of $B$. The purpose of the sign function is to restrict the structure of walks in the bidirected graph: if two edges uvuv and vwvw appear consecutively in a walk, then the sign of edge uvuv at $v$ must be the opposite of the sign of edge vwvw at $v$. Stated informally: every walk must change signs when passing through a vertex.
Bidirected graphs were introduced independently by Kotzig [Kotzig1959, Kotzig1959b, Kotzig1960] and by Edmonds and Johnson [edmonds2003matching]. Recently, much research on bidirected graphs has been motivated by the fact that they are related to the structure of perfect matchings in (undirected) graphs. In particular, there is a way to translate between undirected graphs with perfect matchings and bidirected graphs. Suppose that $G$ is an undirected graph with a perfect matching $M$. We construct a bidirected graph $(B,\sigma)$ from $G$ as follows. For each edge uvuv of $M$:
  • choose an end $u$ arbitrarily;
  • assign every edge uwuw with $w\neq v$ the sign - at $u$;
  • assign every other edge vwvw with $w\neq u$ the sign ++ at $v$; and
  • contract edge uvuv.
Because $M$ is a perfect matching, this procedure assigns every end of every edge a sign. This procedure is reversible: from every bidirected graph, we can construct a corresponding undirected graph with a perfect matching.
There are several open problems and research directions in structural graph theory dedicated to studying the structure of perfect matchings in graphs. Matching minors and perfect matching width were introduced as a substructure and a width parameter that interact well with the structure of matchings in graphs. Norin proposed the Matching Grid Conjecture, an analog of the Grid Minor Theorem, to relate these two concepts. The conjecture states that, given a graph with a perfect matching, either the graph has bounded perfect matching width or it contains a large grid as a matching minor. Using the correspondence between graphs with perfect matchings and bidirected graphs, one can show that matching minors are a bidirected graph analog of butterfly minors (which are themselves a directed graph analog of minors). In fact, by translating matching minors, perfect matching width, and the Matching Grid Conjecture to the bidirected graph setting, the Matching Grid Conjecture can be understood as a Bidirected Grid Conjecture, i.e. a conjectured grid theorem for bidirected graphs.
Recently, Hatzel, Rabinovich, and Wiederrecht [hatzel2019cyclewidth] proved the Matching Grid Conjecture for bipartite graphs. Their proof relied on two important insights. First, if we start with a bipartite graph $G$ with a perfect matching and construct a bidirected graph $(B,\sigma)$ from $G$ using the procedure described above, the bidirected graph $(B,\sigma)$ is in fact a directed graph, i.e. every edge in $(B,\sigma)$ can be traversed in at most one direction. They then used the Directed Grid Theorem to prove the Matching Grid Conjecture in this case.
Inspired by the importance of understanding connectivity to the Matching Grid Conjecture, we study the connectivity of bidirected graphs. Because of the way that walks are defined in bidirected graphs, trails and paths have more complicated structures than in undirected or directed graphs. For example, if a bidirected graph $B$ contains a trail from $x$ to $y$, it may not necessarily contain a path from $x$ to $y$. Since connectivity in bidirected graphs is so complex, we make two assumptions: we work with rooted graphs (and rooted connectivity), and we assume further that the bidirected graphs are clean, meaning that they exclude certain types of trails.
The main result of this paper is a structure theorem for rooted clean bidirected graphs that models any given bidirected graph on a directed graph. The rooted connectivity of the bidirected graph can then be determined through the rooted connectivity of the digraph model. We in fact have two versions of this structure theorem, one tailored to studying edge connectivity and one tailored to vertex connectivity. Using these structure theorems, we prove versions of Lovasz’s flame theorem, Pym’s theorem, and Menger’s theorem for edge-disjoint and vertex-disjoint paths in bidirected graphs. We also give examples showing that these results may fail without our assumption of cleanness.
$1.$ Overview of the proof
The goal of our main result is to model bidirected graphs on directed graphs that represent their connectivity. Let $B$ be a rooted clean bidirected graph. First, we partition the edges of $B$ into two groups: the directable edges, which can be traversed in at most one direction on paths (resp. trails) from the root, and the undirectable edges, which can be traversed in both directions in this way. (If every edge of $B$ is directable, then $B$ is in fact a directed graph.) The set of directable edges forms the basis for the directed graph on which we model $B$. The undirectable edges, meanwhile, can be partitioned into connected components of vertices that each have connectivity one from the root of $B$. These components are represented in the directed graph via a single vertex.
All of our results have algorithmic versions. The main difficulty is to find an algorithm which checks in polynomial time whether an edge is trail-directable. However, as explained in [bowler2023mengertheorembidirected], such an algorithm can be quickly obtained from Edmonds’ Blossom algorithm. Given this, it should be clear from our constructions that each of our decompositions of a bidirected graph can also be constructed in polynomial time, which yields efficient algorithms for all our applications.
1.2. Structure of the paper
We begin by introducing the notation related to bidirected graphs in section 2. In LABEL:sec:_edge_decmp, we prove our main result, a structure theorem for rooted connectivities of bidirected graphs. In LABEL:sec:edge-menger, LABEL:sec:edge-flame and LABEL:sec:edge-pym, we use our structure theorem to generalize the edge-disjoint versions of Menger’s theorem, Lovász flame theorem, and Pym’s theorem, respectively, to edge-clean bidirected graphs. In LABEL:sec:_from_edge_to_vertex, we derive the vertex-disjoint variants of these applications for bidirected graphs. Finally, we adapt our structure theorem to the vertex-disjoint context in LABEL:sec:_vertex-decomp.
2. Preliminaries and Notation
For the standard notation for undirected graphs we refer the reader to [diestel2024graph]. A bidirected graph $B=(G,\sigma)$ consists of a loopless multigraph $G=(V,E)$, a corresponding set of half-edges defined as
$\textbf{E}(B):=\left\{(e,u)\colon e\in E\text{ and }u\text{ is an endpoint of }e\right\}$
and a signing $\sigma:\textbf{E}\to\left\{+,-\right\}$ assigning to each half-edge $(e,u)$ its sign $\sigma(e,u):=\sigma((e,u))$; we say that $e$ has sign $\sigma(e,u)$ at $u$. Then $V(B):=V$ is the vertex set of $B$ and $E(B):=E$ is its edge set. We refer to the elements of $V(B)$ and $E(B)$ as the vertices and the edges of $B$, respectively. A signed vertex of $B$ is a pair $(v,\alpha)$ of a vertex $v$ of $B$ and a sign $\alpha\in\{+,-\}$. Given a vertex set $X$, we let $\mathcal{V}(X):=\{(v,\alpha):\ v\in V(B),\alpha\in\{+,-\}\}$. For an edge set $F$, we write $V(F)$ for the set of the endpoints of the edges in $F$.
An oriented edge

### v2

@@
 In this paper, we show a structure theorem about rooted connectivity in bidirected graphs in terms of directed graphs.
-As applications, we prove Lovász’ flame theorem, Pym’s theorem and a strong variant of Menger’s theorem for a class of bidirected graphs and provide counterexamples in the general case.
+As applications, we prove Lovász’s flame theorem, Pym’s theorem and a strong variant of Menger’s theorem for a class of bidirected graphs and provide counterexamples in the general case.
 Key words and phrases:
@@
 A bidirected graph is a graph $B$ endowed with a sign function $\sigma$ that assigns a sign $\{+,-\}$ to each end of every edge of $B$.
-The purpose of the sign function is to restrict the structure of walks in the bidirected graph: if two edges uvuv and vwvw appear consecutively in a walk, then the sign of edge uvuv at $v$ must be the opposite of the sign of edge vwvw at $v$.
+The purpose of the sign function is to restrict the structure of walks in the bidirected graph: if two edges $uv$ and $vw$ appear consecutively in a walk, then the sign of edge $uv$ at $v$ must be the opposite of the sign of edge $vw$ at $v$.
 Stated informally: every walk must change signs when passing through a vertex.
 Bidirected graphs were introduced independently by Kotzig [Kotzig1959, Kotzig1959b, Kotzig1960] and by Edmonds and Johnson [edmonds2003matching].
-Recently, much research on bidirected graphs has been motivated by the fact that they are related to the structure of perfect matchings in (undirected) graphs.
+Recently, much research on bidirected graphs has been motivated by the fact that they are related to the structure of perfect matchings in (undirected) graphs [wiederrecht2020thesis, bowler2023mengertheorembidirected, bowler2024hitting, bowler2025ConnectoidsI, KITA2017, Nickel2025, Ghorbani2025].
 In particular, there is a way to translate between undirected graphs with perfect matchings and bidirected graphs.
@@
 We construct a bidirected graph $(B,\sigma)$ from $G$ as follows.
-For each edge uvuv of $M$:
+For each edge $uv$ of $M$:
   • choose an end $u$ arbitrarily;
-  • assign every edge uwuw with $w\neq v$ the sign - at $u$;
-  • assign every other edge vwvw with $w\neq u$ the sign ++ at $v$; and
-  • contract edge uvuv.
+  • assign every edge $uw$ with $w\neq v$ the sign - at $u$;
+  • assign every other edge $vw$ with $w\neq u$ the sign ++ at $v$; and
+  • contract edge $uv$.
 Because $M$ is a perfect matching, this procedure assigns every end of every edge a sign.
@@
 Matching minors and perfect matching width were introduced as a substructure and a width parameter that interact well with the structure of matchings in graphs.
-Norin proposed the Matching Grid Conjecture, an analog of the Grid Minor Theorem, to relate these two concepts.
+Norin [norine2005matching] proposed the Matching Grid Conjecture, an analog of the Grid Minor Theorem, to relate these two concepts.
 The conjecture states that, given a graph with a perfect matching, either the graph has bounded perfect matching width or it contains a large grid as a matching minor.
@@
 every edge in $(B,\sigma)$ can be traversed in at most one direction.
-They then used the Directed Grid Theorem to prove the Matching Grid Conjecture in this case.
+They then used the Directed Grid Theorem [kawarabayashi2015directed] to prove the Matching Grid Conjecture in this case.
 Inspired by the importance of understanding connectivity to the Matching Grid Conjecture, we study the connectivity of bidirected graphs.
@@
 We also give examples showing that these results may fail without our assumption of cleanness.
-$1.$ Overview of the proof
+1.1. Overview of the proof
 The goal of our main result is to model bidirected graphs on directed graphs that represent their connectivity.